# A cutting-surface consensus approach for distributed robust optimization of multi-agent systems

### v2

@@ Abstract
A novel and fully distributed optimization method is proposed for the distributed robust convex program (DRCP) over a time-varying unbalanced directed network under the uniformly jointly strongly connected (UJSC) assumption. Firstly, a tractable approximated DRCP (ADRCP) is introduced by discretizing the semi-infinite constraints into a finite number of inequality constraints and restricting the right-hand side of the constraints with a positive parameter. This problem is iteratively solved by a distributed projected gradient algorithm proposed in this paper, which is based on epigraphic reformulation and subgradient projected algorithms. Secondly, a cutting-surface consensus approach is proposed for locating an approximately optimal consensus solution of the DRCP with guaranteed feasibility. This approach is based on iteratively approximating the DRCP by successively reducing the restriction parameter of the right-hand constraints and populating the cutting-surfaces into the existing finite set of constraints. Thirdly, to ensure finite-time termination of the distributed optimization, a distributed termination algorithm is developed based on consensus and zeroth-order stopping conditions under UJSC graphs. Fourthly, it is proved that the cutting-surface consensus approach terminates finitely and yields a feasible and approximate optimal solution for each agent. Finally, the effectiveness of the approach is illustrated through a numerical example.
@@
Distributed robust convex program, time-varying unbalanced directed graphs, semi-infinite constraints, finite-time termination
@@ 1 Introduction
Distributed optimization refers to an optimization strategy for multi-agent systems in which agents agree on a consensus and feasible decision that guarantees the optimality of the global objective through local computation and agent communication. Distributed optimization has been extensively studied [1, 2, 3] and has been applicated in a wide range of fields such as multi-robot system scheduling [4], unmanned aerial vehicle (UAV) formation flight [5], cooperative localization in sensor networks [6], network resource allocation [7], and economic dispatch of smart grids [8]. Uncertain distributed optimization is considerably important, e.g., to guarantee stability or to ensure safety [9]. Uncertainties considered in distributed optimization include inaccurate gradient information [10, 11], communication noise [12, 13], communication time delay [14], data packet loss [15], constraints affected by uncertainty, etc. This article focuses on distributed optimization problems with uncertain constraints.
@@
Techniques for dealing with distributed optimization with uncertain constraints have been widely researched. In the early work, Lee and Nedic [16, 17] introduced distributed random projection algorithms by designing random projections to ensure almost sure feasibility of the solution and using a distributed subgradient algorithm [18] to cooperatively minimize the global objective. In [19, 20, 21, 22, 23, 24, 25], some scenario-based algorithms were proposed, which are based on extracting a large number of scenarios from the uncertainty set to obtain a scenario problem (SP), and then applying existing constrained distributed optimization methods to solve the SP. These algorithms provide probabilistic feasibility of the resulting solutions. The above two algorithms require that the uncertainty parameters be randomly sampled from an underlying distribution. Thus, they unavoidably fail to find a feasible solution for all values of the uncertainty but can provide feasibility assurance in a probabilistic sense. However, in some practical applications, all agents need to locate guaranteed feasible and optimal consensus solutions for a given range of uncertainty, regardless of the distribution, to ensure that the multi-agent system has a good robustness performance.
@@
To address such issues, this paper considers a distributed robust convex program (DRCP) where the inequality constraints are influenced by bounded uncertainty. According to the treatment of uncertainty, the existing methods are categorized into two main types: the robust counterpart method [26, 27, 28] and the cutting-plane consensus algorithm [29, 30]. The former presents a direct strategy by transforming the DRCP into an equivalent and computationally tractable problem based on duality theory, which is then solved by an existing distributed algorithm [26]. This method asymptotically converges to a feasible optimal solution, whereas it is primarily used for uncertain linear and conic quadratic programs (see the preface of [9]). Moreover, this method requires the global information of constraints and uncertainty sets, and thus, the system considered by this method is not fully distributed. The later is a fully distributed algorithm that relies on iteratively extending cutting-planes into the finite set of constraints to approximate the DRCP, thereby enabling the algorithm to asymptotically converge to the optimal solution of the DRCP. However, this polyhedral approximation framework results in an outer approximation [29], which does not guarantee the feasibility of the solution. To the best of our knowledge, none of the existing fully distributed algorithms guarantee the feasibility of the solution in the DRCP. Besides, these algorithms for solving the DRCP fail to achieve finite-time termination, which is essential for practice. Therefore, the motivation of this article is to propose a novel and fully distributed approach for finding a feasible and approximate optimal consensus solution for the DRCP within a finite number of iterations.
@@
From the aforementioned motivation, a cutting-surface consensus approach is proposed based on the centralized cutting-surface method [31] and the right-hand restriction approach [32, 33] for solving the DRCP. Firstly, an approximated DRCP (ADRCP) is constructed by discretizing the semi-infinite constraints into a finite number of inequality constraints and restricting the right-hand side of the constraints with a positive parameter to retrieve computational tractability. Secondly, the ADRCP is solved by the distributed projected gradient algorithm, which is based on epigraphic reformulation [34] and the subgradient projected algorithm [35], and the resulting optimal solution is used to approximate that of the DRCP. Thirdly, a lower level program (LLP) is introduced to determine whether the obtained solution of the ADRCP is feasible to the original problem (DRCP). If infeasible, the ADRCP is tightened by populating the uncertain parameter corresponding to the maximum constraint violation into the existing finite set; otherwise, a feasible point of the DRCP is located and used to update the optimal solution candidate. Finally, if the stopping criterion is satisfied, the optimal solution candidate is taken as the desired approximate optimal solution of the DRCP; otherwise, the restriction parameters is reduced and then the routine is performed again with updated information based on the explored solution of the ADRCP until the desired solution is located. Based on the above procedures, a cutting-surface consensus approach has been formulated to converge to an optimal solution of the DRCP. The main contributions of this paper are summarized as follows.
@@
    The DRCP with nonidentical constraints and nonidentical uncertainty sets is considered over a time-varying unbalanced directed network under the uniformly jointly strongly connected (UJSC) assumption since such a problem is more favorable in practical applications.
@@
    A novel and fully distributed cutting-surface consensus approach is proposed for solving the DRCP. Unlike the above-mentioned methods in [26, 27, 28, 30, 29], the proposed approach terminates finitely and yields a feasible solution for each agent satisfying consensus and zeroth-order stopping conditions to certain tolerances.
@@
    A distributed projected gradient (DPG) algorithm is developed for solving the ADRCP problem. Besides, we prove its convergence and convergence rate of $O({\ln t}/{\sqrt{t}})$. Unlike the D-RFP algorithm in [34] that reaches almost sure feasibility of the solution, the DPG algorithm adopts Euclidean projection that ensures the feasibility of the solution to prevent infinite iterations of the cutting-surface consensus approach.
@@
    Based on the zeroth-order stopping conditions [36] and finite-time consensus algorithm [37], a distributed termination algorithm is developed and applied to the process of solving the ADRCP and the DRCP. Unlike the finite-time consensus technique [38, 39, 37] that only guarantees approximate consensus of the resulting solutions under strongly connected graphs, our algorithm can be used in UJSC networks and introduces zeroth-order stopping conditions to ensure all agents obtain approximate optimal solutions for convergent distributed optimization algorithms in the discrete-time setting.
  5. 5.
    It is mathematically proven that the cutting-surface consensus approach terminates within finite iterations.
The remaining parts of the paper are organized as follows. Section 2 presents the distributed robust convex optimization problem and network topology. Section 3 states the mathematical description and stopping criterion of the ADRCP and presents the associated algorithms. Section 4 presents the cutting-surface consensus approach for locating a feasible solution of the DRCP that reaches consensus and zeroth-order stopping conditions to certain tolerances within finite iterations, and it also presents a mathematical proof of this result. Section 5 analyzes the effect of the algorithm with a numerical example and compares it with existing methods. Section 6 provides conclusions and an outlook on future work.
Notation: $\mathbb{R}^{m}$ denotes the set of all $m$ dimensional real column vector; $\mathbb{Z}$ denotes the set of all nonnegative integers; $x^{\top}$ and $A^{\top}$ denote the transpose of the vector $x$ and the matrix $A$; $I$ represents an identity matrix with compatible size; ${\bf 1}_{m}$ denotes the $m$-dimensional column vector with all entries equal to 1; $e_{i}$ is a column vector where the $i$-th element equals to 1 and the remaining elements equal to 0; $\|x\|$ represents the Euclidean norm of the vector $x$; $|a|$ denotes the absolute value of a scalar $a\in\mathbb{R}$; $\rm{dist}(\overline{x},X)$ denotes the Euclidean distance from a vector $\overline{x}$ to a set $X$, i.e. $\rm{dist}(\overline{x},X)={\inf}_{x\in X}\|\overline{x}-x\|$; $P_{X}[\overline{x}]$ denotes the Euclidean projection of a vector $\overline{x}$ on a closed convex set $X$, i.e. $P_{X}[\overline{x}]=\rm{arg}\min_{x\in X}\|\overline{x}-x\|$.
@@ 2 Problem Statement
$&\mathop{\min}\limits_{{x}\in X}&&F(x)=\sum^{m}_{i=1}f_{i}({x})\\ &\rm{s.t.}&&g_{i}({x},y_{i})\leq 0\quad\forall y_{i}\in Y_{i}\quad i=1,2,...m\end{aligned}$ (${\rm DRCP}$)
where ${x}$ is a decision vector of agents, $y_{i}$ is an uncertainty parameter of agent $i$, the global constraint set $X\subseteq\mathbb{R}^{n}$ is a non-empty and compact convex set, $Y_{i}\subseteq\mathbb{R}$ is a nonempty and compact convex set for all $i\in\mathcal{V}$, $f_{i}({x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the local cost function of agent $i$, $F({x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$ is the global objective function, $g_{i}({x},y_{i}):\mathbb{R}^{n}\times Y_{i}\rightarrow\mathbb{R}$ is the local constraint of agent $i$. Both $f_{i}$ and $g_{i}$ are assumed to be convex functions with respect to ${x}$. Besides, $g_{i}$ is assumed to be continuous on $X\times Y_{i}$ for all $i\in\mathcal{V}$. We set $x^{*}$ as an optimal solution of DRCP, and $x^{*}$ belongs to the optimal solution set $\mathcal{F}^{*}$.
Each agent only knows its local cost function $f_{i}$, constraints $g_{i}$, and uncertainty set $Y_{i}$, and this information can be different for each agent. It is defined that the local feasible set of agent $i$ is $\mathcal{F}_{i}=\left\{x\in X|g_{i}(x,y_{i})\leq 0,\forall y_{i}\in Y_{i}\right\}$. To ensure the solvability of the DRCP, we make the following assumptions.
Assumption 1 (Solvability and Interior Point)
$\mathcal{F}$ is not an empty set, and there is a point $\hat{x}\in\mathcal{F}$ satisfying
$g_{i}(\hat{x},y_{i})<0,\forall y_{i}\in Y_{i}\quad i=1,...,m$
where $\mathcal{F}$ is the intersection of local feasible sets of all agents, i.e. $\mathcal{F}=\cap^{m}_{i=1}\mathcal{F}_{i}$.
@@
[…]
The communication among agents is represented by time-varying directed graphs $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t))$, where $\mathcal{V}=\left\{1,...,m\right\}$ denotes the vertex set, and $\mathcal{E}(t)=\left\{1,...,m\right\}^{2}$ is the edge set in the time slot $t\in\mathbb{Z}$. A directed edge $(i,j)\in\mathcal{E}(t)$ indicates that agent $i$ can directly transmit information to agent $j$ in the time slot $t$. $(i,i)$ represents the self-loop of agent $i$. $\mathcal{G}(t)$ is strongly connected if every agent is reachable by all other agents. For any agent $i,j\in\mathcal{V}$, the minimum number of edges between two agents is called the distance from $i$ to $j$, denoted as ${\rm dist}(i,j)$. The diameter of the graph is denoted as ${\rm diam}(\mathcal{G}(t))=\mathop{\max}\limits_{\forall i,j\in\mathcal{V}}\ {\rm dist}(i,j)$. In the time slot $t$, for agent $i$, its out-neighbors set is $N_{i}^{out}(t)=\left\{j|(i,j)\in\mathcal{E}(t)\right\}$, and its in-neighbors set is $N_{i}^{in}(t)=\left\{j|(j,i)\in\mathcal{E}(t)\right\}$. Define the weight matrix $A(t)=[a_{ij}(t)]\in\mathbb{R}^{m\times m}$, which satisfies that $a_{ij}(t)>0$ if $j\in N_{i}^{in}(t)$ and $a_{ij}(t)=0$, otherwise. If $\sum^{m}_{i=1}a_{ij}(t)=1$ for any $j\in\mathcal{V}$, then the matrix $A(t)$ is column stochastic. If $\sum^{m}_{j=1}a_{ij}(t)=1$ for any $i\in\mathcal{V}$, then the matrix $A(t)$ is row stochastic. The matrix $A(t)$ is doubly stochastic if it satisfies both the column and row stochastic conditions. In order to ensure that the information from any agent can be directly or indirectly transmitted to all other agents within finite time slots, the following assumption is made.
@@
The time-varying graphs $\left\{\mathcal{G}(t)\right\}$ are uniformly jointly strongly connected (UJSC), i.e. for all $t\geq 0$, there is an interger $S>0$ making $\mathcal{G}(t:t+S)$ strongly connected.
$\mathcal{G}(t:t+S)=([m],\mathcal{G}(t)\cup...\cup\mathcal{G}(t+S-1)$
The objective of our work is to propose an approach to solve the (${\rm DRCP}$) within a finite number of iterations under time-varying unbalanced directed graphs while guaranteeing the feasibility of the approximate optimal solution.
3 Approximation Problem and Finite-Time Termination Method
This section introduces an approximation problem of the (${\rm DRCP}$) and proposes a distributed projected gradient (DPG) algorithm for solving this problem, which can be used in time-varying unbalanced directed networks under UJSC assumption. Then, we present the convergence of the DPG algorithm and discuss its convergence rate. Besides, we develop a distributed finite-time termination algorithm. By combining it with the DPG algorithm, all agents obtain approximate optimal and consensus solutions of the ADRCP within finite time slots.
3.1 Approximation Problem
Motivated by the upper bounding problem in [32], an approximation problem of the (${\rm DRCP}$) is introduced by discretizing the compact sets $Y_{i}$ into the finite sets $Y^{k}_{i}$ and restricting the right-hand sides of the constraints with positive parameters $\epsilon^{k}_{i}$. The mathematical description of the approximation problem is as follows.
$&\mathop{\min}\limits_{{x}\in X}&&F(x)\!=\!\sum^{m}_{i=1}f_{i}({x})\\ &\rm{s.t.}&&g_{i}({x},y_{i})\!\leq\!-\epsilon^{k}_{i}\ \ \forall y_{i}\!\in\!Y^{k}_{i}\ \ i=1,2,...,m\end{aligned}$ ($\rm{ADRCP}^{k}$)
where $\epsilon^{k}_{i}$ is a restriction parameter of agent $i$, and $\epsilon^{k}_{i}>0$, $Y^{k}_{i}$ is a finite set of agent $i$, and $Y^{k}_{i}\subset Y_{i}$. We set $x^{k,*}$ as an optimal solution of the ($\rm{ADRCP}^{k}$), and $x^{k,*}$ belongs to the optimal solution set $\mathcal{F}^{k,*}$.

### v3

@@ Abstract
A novel and fully distributed optimization method is proposed for the distributed robust convex program (DRCP) over a time-varying unbalanced directed network under the uniformly jointly strongly connected (UJSC) assumption. Firstly, an approximated DRCP (ADRCP) is introduced by discretizing the semi-infinite constraints into a finite number of inequality constraints to ensure tractability and restricting the right-hand side of the constraints with a positive parameter to ensure a feasible solution for (DRCP) can be obtained. This problem is iteratively solved by a distributed projected gradient algorithm proposed in this paper, which is based on epigraphic reformulation and gradient projected operations. Secondly, a cutting-surface consensus approach is proposed for locating an approximately optimal consensus solution of the DRCP with guaranteed local feasibility for each agent. This approach is based on iteratively approximating the DRCP by successively reducing the restriction parameter of the right-hand constraints and adding the cutting-surfaces into the existing finite set of constraints. Thirdly, to ensure finite-time termination of the distributed optimization, a distributed termination algorithm is developed based on consensus and zeroth-order stopping conditions under UJSC graphs. Fourthly, it is proved that the cutting-surface consensus approach terminates finitely and yields a feasible and approximate optimal solution for each agent. Finally, the effectiveness of the approach is illustrated through a numerical example.
@@
Distributed robust convex program, time-varying unbalanced directed graphs, semi-infinite constraints, local feasibility, finite-time termination
@@ 1 Introduction
Distributed optimization refers to an optimization strategy for multi-agent systems in which agents agree on a consensus and feasible decision that guarantees the optimality of a global objective through local computation and agent communication. Distributed optimization has been extensively studied [1, 2, 3] and has been applied in a wide range of fields such as multi-robot system scheduling [4], unmanned aerial vehicle (UAV) formation flight [5], cooperative localization in sensor networks [6], network resource allocation [7], and economic dispatch of smart grids [8]. Uncertain distributed optimization is considerably important, e.g., to guarantee stability or to ensure safety [9]. Uncertainties considered in distributed optimization include inaccurate gradient information [10], communication noise [11], communication time delay [12], data packet loss [13], constraints affected by uncertainty, etc. This article focuses on distributed optimization problems with uncertain constraints.
@@
Techniques for dealing with distributed optimization with uncertain constraints have been widely researched. In the early work, Lee and Nedic [14, 15] introduced distributed random projection algorithms by designing random projections to ensure almost sure feasibility of the solution and using a distributed subgradient algorithm [16] to cooperatively minimize the global objective. In [17, 18, 19, 20, 21, 22], some scenario-based algorithms were proposed, which are based on extracting a large number of scenarios from the uncertainty set to obtain a scenario problem (SP), and then applying existing constrained distributed optimization methods to solve the SP. These algorithms provide probabilistic feasibility of the resulting solutions. The above two algorithms require the realizations of the uncertainty parameters to be randomly sampled from an underlying distribution. Thus, they unavoidably fail to find a feasible solution for all values of the uncertainty but can provide feasibility assurance in a probabilistic sense. However, in some practical applications, uncertain constraints are safety-critical, and hence, all agents need to locate guaranteed feasible and optimal consensus solutions to ensure that the multi-agent system has a good robustness performance.
@@
To address such issues, this paper considers a distributed robust convex program (DRCP) where the inequality constraints are influenced by bounded uncertainty. According to the treatment of uncertainty, the existing methods are categorized into two main types: the robust counterpart method [23, 24, 25] and the cutting-plane consensus algorithm [26, 27]. The former presents a direct strategy by transforming the DRCP into an equivalent and computationally tractable problem based on duality theory, which is then solved by an existing distributed algorithm [23]. This method asymptotically converges to a feasible optimal solution, whereas it is primarily used for uncertain linear and conic quadratic programs (see the preface of [9]). Moreover, this method requires the global information of constraints and uncertainty sets, and thus, the system considered by this method is not fully distributed. The later is a fully distributed algorithm that relies on iteratively extending linear cutting-planes into the finite set of constraints to approximate the DRCP, thereby enabling the algorithm to asymptotically converge to the optimal solution of the DRCP. However, this polyhedral approximation framework for the feasible set of the DRCP results in an outer approximation [26], which does not guarantee the feasibility of the solution. To the best of our knowledge, none of the existing fully distributed algorithms guarantee the feasibility of the solution in the DRCP. Additionally, these algorithms for solving the DRCP fail to achieve finite-time termination, which is essential for practice. Therefore, the motivation of this article is to propose a novel and fully distributed approach for finding a feasible and approximate optimal consensus solution for the DRCP within a finite number of iterations.
@@
From the aforementioned motivation, a novel cutting-surface consensus approach is proposed to solve the DRCP. The method is based on the constraint-handling strategies of the centralized cutting-surface algorithm [28], and the centralized right-hand restriction algorithm [29] which is used to deal with standard semi-infinite programs. Firstly, an approximated DRCP (ADRCP) is constructed by discretizing the semi-infinite constraints into a finite number of inequality constraints to retrieve computational tractability and restricting the right-hand side of the constraints with a positive parameter to ensure that a feasible solution for (DRCP) can be attained. Secondly, the ADRCP is solved by the distributed projected gradient algorithm, which is based on epigraphic reformulation [30] and the gradient projected algorithm [31], and the resulting optimal solution is used to approximate that of the DRCP. Thirdly, a lower level program (LLP) is introduced to determine whether the obtained solution of the ADRCP is feasible to the original problem (DRCP). If infeasible, the ADRCP is tightened by adding a convex and possibly nonlinear cutting-surface based on the maximum constraint violation into the existing finite constraint set; otherwise, a feasible point of the DRCP is located and used to update the optimal solution candidate. Finally, if the stopping criterion is satisfied, the optimal solution candidate is taken as the desired approximate optimal solution of the DRCP; otherwise, the restriction parameters is reduced and then the routine is performed again with updated information based on the explored solution of the ADRCP until the desired solution is located. Based on the above procedures, a cutting-surface consensus approach has been formulated to converge to an optimal solution of the DRCP. The main contributions of this paper are summarized as follows.
@@
    The DRCP with nonidentical constraints and nonidentical uncertainty sets is considered over a time-varying unbalanced directed network under the uniformly jointly strongly connected (UJSC) assumption since such a problem is more favorable in practical applications. To solve this problem, this paper proposes a novel and fully distributed cutting-surface consensus approach. Unlike the above-mentioned methods in [23, 24, 25, 27, 26], the proposed approach terminates finitely and yields a locally feasible solution for each agent satisfying consensus and zeroth-order stopping conditions to certain tolerances.
@@
    A distributed projected gradient (DPG) algorithm is developed for solving the ADRCP problem. In addition, we prove its convergence and convergence rate of $O({\ln t}/{\sqrt{t}})$. Compared to existing subgradient projection algorithms [31, 32, 33, 34, 35, 36], the proposed DPG algorithm can be applied to the distributed constrained optimization problems with nonidentical constraints under time-varying unbalanced directed network.
@@
    Based on the zeroth-order stopping conditions [37] and finite-time consensus algorithm [38], a distributed termination algorithm is developed and applied to the process of solving the ADRCP and the DRCP. Unlike the finite-time consensus technique [38, 39, 40] that only guarantees approximate consensus of the resulting solutions under strongly connected graphs, our algorithm can be used in UJSC networks and introduces zeroth-order stopping conditions to prevent the convergent distributed optimization algorithms from terminating prematurely on a very steep slope or a flat plateau. This ensures that all agents obtain approximate optimal solutions in the discrete-time setting.
@@
    It is mathematically proven that the cutting-surface consensus approach reaches termination in finite iterations and guarantees local feasibility of the resulting solutions.
The remaining parts of the paper are organized as follows. Section 2 presents the distributed robust convex optimization problem and network topology. Section 3 presents our main result of locating a feasible solution of the DRCP that reaches consensus and zeroth-order stopping conditions to certain tolerances within finite iterations, and it also presents a mathematical proof of this result. Section 4 analyzes the effect of the algorithm with a numerical example and compares it with existing methods. Section 5 provides conclusions and an outlook on future work.
Notation: $\mathbb{R}^{m}$ denotes the set of all $m$ dimensional real column vector; $\mathbb{Z}$ denotes the set of all nonnegative integers; $x^{\top}$ and $A^{\top}$ denote the transpose of the vector $x$ and the matrix $A$; II represents an identity matrix with compatible size; ${\bf 1}_{m}$ denotes the $m$-dimensional column vector with all entries equal to 1; $e_{i}$ is a column vector where the $i$-th element equals to 1 and the remaining elements equal to 0; $\|x\|$ represents the Euclidean norm of the vector $x$; $|a|$ denotes the absolute value of a scalar $a\in\mathbb{R}$; ${\rm{dist}}(\overline{x},X)$ denotes the Euclidean distance from a vector $\overline{x}$ to a set $X$, i.e. ${\rm{dist}}(\overline{x},X)={\inf}_{x\in X}\|\overline{x}-x\|$; $P_{X}[\overline{x}]$ denotes the Euclidean projection of a vector $\overline{x}$ on a closed convex set $X$, i.e. $P_{X}[\overline{x}]={\rm{arg}}\min_{x\in X}\|\overline{x}-x\|$.
@@ 2 Problem Statement
$\begin{aligned} &\mathop{\min}\limits_{{x}\in X}&&F(x)=\sum^{m}_{i=1}f_{i}({x})\\ &\ \rm{s.t.}&&g_{i}({x},y_{i})\leq 0\quad\forall y_{i}\in Y_{i}\quad i=1,2,\ldots,m,\end{aligned}$ ($\rm DRCP$})
where x{x} is a common decision vector, $y_{i}$ is an uncertainty parameter of agent $i$, the global constraint set $X\subseteq\mathbb{R}^{n}$ is a nonempty and compact convex set, $Y_{i}\subseteq\mathbb{R}$ is a nonempty and compact convex set for any $i\in\{1,\ldots,m\}$, $f_{i}({x}):X\rightarrow\mathbb{R}$ is the local cost function of agent $i$, $F({x}):X\rightarrow\mathbb{R}$ is the global objective function, $g_{i}({x},y_{i}):X\times Y_{i}\rightarrow\mathbb{R}$ is the local constraint function of agent $i$. Assume $f_{i}$ and $g_{i}$ are convex with respect to x{x}. Additionally, $g_{i}$ is assumed to be continuous on $X\times Y_{i}$ for any $i\!\in\!\{1,\ldots,m\}$. Set $x^{*}$ as an optimal solution of DRCP, and $x^{*}$ belongs to the optimal solution set $\mathcal{F}^{*}$.
Each agent only knows its local cost function $f_{i}$, constraint $g_{i}$, and uncertainty set $Y_{i}$, and this information can be different for each agent. It is defined that the local feasible set of agent $i$ is $\mathcal{F}_{i}=\left\{x\in X|g_{i}(x,y_{i})\leq 0,\forall y_{i}\in Y_{i}\right\}$. To ensure the feasibility of (DRCP), we make the following assumption.
Assumption 1 (Feasibility of $\rm(DRCP)$} and Slater Point)
$\mathcal{F}$ is a nonempty set, and there is a point $\hat{x}\in\mathcal{F}$ satisfying
$g_{i}(\hat{x},y_{i})<0,\ \ \forall y_{i}\in Y_{i},\ \ \forall i\in\{1,\ldots,m\},$
where $\mathcal{F}$ is the intersection of local feasible sets of all agents, i.e., $\mathcal{F}=\bigcap^{m}_{i=1}\mathcal{F}_{i}$.
@@
Since an optimization problem with uncertain local cost functions can be transformed into an epigraph reformulation problem with the same form as the ($\rm DRCP$}), it is without loss of generality that this paper only considers the uncertainty appearing in the local constraints.
The communication among agents is represented by time-varying directed graphs $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t))$, where $\mathcal{V}=\left\{1,\ldots,m\right\}$ denotes the vertex set, and $\mathcal{E}(t)=\left\{1,\ldots,m\right\}^{2}$ is the edge set at time $t\in\mathbb{Z}$. A directed edge $(i,j)\in\mathcal{E}(t)$ indicates that agent $i$ can directly transmit information to agent $j$ at time $t$. $(i,i)$ represents the self-loop of agent $i$. $\mathcal{G}(t)$ is strongly connected if every agent is reachable by all other agents. For any pair of agents $(i,j)$ with $i,j\in\mathcal{V}$, the minimum number of edges between two agents is called the distance from $i$ to $j$, denoted as ${\rm dist}(i,j)$. The diameter of the graph is denoted as ${\rm diam}(\mathcal{G}(t))=\mathop{\max}\limits_{\forall i,j\in\mathcal{V}}\ {\rm dist}(i,j)$. At the time $t$, for agent $i$, its out-neighbors set is $N_{i}^{out}(t)=\left\{j|(i,j)\in\mathcal{E}(t)\right\}$, and its in-neighbors set is $N_{i}^{in}(t)=\left\{j|(j,i)\in\mathcal{E}(t)\right\}$. Define the weight matrix $A(t)=[a_{ij}(t)]\in\mathbb{R}^{m\times m}$, which satisfies that $a_{ij}(t)>0$ if $j\in N_{i}^{in}(t)$ and $a_{ij}(t)=0$, otherwise. In order to ensure that the information from any agent can be directly or indirectly transmitted to all other agents in finite time, the following assumption is made.
@@
The time-varying graphs $\left\{\mathcal{G}(t)\right\}$ are uniformly jointly strongly connected (UJSC), i.e. for all $t\geq 0$, there is an integer $S>0$ making $\mathcal{G}(t:t+S)$ strongly connected, where
$\mathcal{G}(t:t+S)=(\mathcal{V},\mathcal{E}(t)\cup\ldots\cup\mathcal{E}(t+S-1)).$
The objective of our work is to propose an approach to solve the ($\rm DRCP$}) within a finite number of iterations under time-varying unbalanced directed graphs while guaranteeing the local feasibility of the approximate optimal solution.
3 Main results
Since the constraints of the ($\rm DRCP$}) are semi-infinite constraints, which means that an infinite number of inequality constraints are imposed on finite-dimensional decision variables, solving this problem is NP-hard. To solve the ($\rm DRCP$}), this section firstly introduces an approximation problem of the ($\rm DRCP$}) and proposes a distributed projected gradient (DPG) algorithm for solving this approximation problem under the setting of time-varying unbalanced directed networks. Then, we develop a distributed finite-time termination algorithm. By combining it with the DPG algorithm, all agents obtain approximate optimal and consensus solutions of the approximation problem within finite time. Next, we propose a distributed cutting-surface consensus algorithm that, by tuning the parameters of the approximation problem in each iteration, enables each agent to attain a locally feasible solution of the ($\rm DRCP$}) satisfying consensus and zeroth-order stopping conditions to certain tolerances within a finite number of iterations. Finally, the main results on finite-time termination and local feasibility of the resulting solutions for the distributed cutting-surface consensus algorithm are established.
3.1 Approximation Problem of DRCP and Solving Method
Motivated by the upper bounding problem in [29], an approximation problem of the ($\rm DRCP$}) is introduced by discretizing the compact sets $Y_{i}$ into the finite sets $Y^{k}_{i}$ and restricting the right-hand sides of the constraints with positive parameters $\epsilon^{k}_{i}$ for all $i\in\mathcal{V}$. The mathematical description of the approximation problem is as follows.
$\hskip-5.69054pt\begin{aligned} &\mathop{\min}\limits_{{x}\in X}&&F(x)\!=\!\sum^{m}_{i=1}f_{i}({x})\\ &\ \rm{s.t.}&&g_{i}({x},y_{i})\!\leq\!-\epsilon^{k}_{i}\ \ \forall y_{i}\!\in\!Y^{k}_{i}\ \ i=1,\ldots,m,\end{aligned}$ ($\rm{ADRCP}^{\it k}$)
where for all $i\in\mathcal{V}$, $\epsilon^{k}_{i}$ is a restriction parameter of agent $i$, $\epsilon^{k}_{i}>0$, and $Y^{k}_{i}$ is a finite subset of $Y_{i}$, $Y^{k}_{i}\subset Y_{i}$. We set $x^{k,*}$ as an optimal solution of the ($\rm{ADRCP}^{\it k}$), and $x^{k,*}$ belongs to the optimal solution set $\mathcal{F}^{k,*}$.